3.4 0,00030.0003 0.0003 0.0003 00003 0.00030.0003 0.0003 0.0003 0.0002 3.3 0,0005 0.0005 0.0005 0,0004 0.0004 0.0004 0.0004 0.0004 0,0004 0.0003 3.2 0,00070,0007 0,0006 0.0006 0.0006 0.0006 0,0006 0,0005 0.0005 0.0005 -3. 0,0010 0,0009 0.0009 0,0009 0,0008 0.0008 0,0008 0,0008 0.0007 0,0007 3.0 0,0013 0.0013 0.0013 0.0012 0.0012 0.0011 0.0011 0.0011 0,0010 0.0010 -2.9 0,0019 0.0018 0.0018 0.0017 0.0016 0.0016 0.0015 0,0015 0.0014 0.0014 -2.8 0,00260.0025 0.0024 0.0023 00023 0.0022 0.002 0,0021 0.0020 0.0019 2.7 0,0035 0,0034 0.0033 0.0032 0.0031 0.0030 0.0029 0,00280.0027 0.0026 -2.6 0.0047 0.0045 00044 0.0043 0.004 0.0040 0.0039 0,0038 0,0037 0.0036 -2.5 0,0062 0.0060 0.0059 0.0057 0.0055 0.004 0.0052 0,0051 0.0049 0.0048 24 0,0082 0,0080 0.00780.0075 0.0073 0.0071 0.0069 0,0068 0.0066 0.0064 2.3 0,0107 0.0104 00102 0.0099 0.0096 00094 0.009 0,0089 0.0087 0.0084 -22 0.0139 0.0136 00132 00129 00125 0.0122 0.0119 0.0116 0.0113 0.0110 -2. 0,0179 0.0174 0.0170 00166 0.0162 0.0158 00154 0.0150 0.0146 0.0143 -2.0 0.02280.0222 0.0217 0.0212 0.0207 0.0202 0.0197 0.0192 0.0188 0.0183 -1.9 0.02870.028 0.0274 0.0268 0.0262 0.0256 0.0250 0.0244 0.0239 0.0233 -1.8 0,0359 0035 00344 0.0336 00329 0.0322 0.0314 0.0307 0.0301 0.0294 1.7 0.0446 0.0436 0.0427 0.0418 0.0409 0.040 0.0392 0.0384 0.0375 0.0367 -16 0.054800537 0.0526 0.0516 00505 0.0495 0.0485 0.0475 0.0465 0.0455 15 0.0668 0.0655 0.0643 0.0630 0.0618 0.0606 00594 0.0582 0.0571 0.0559 14 0,0808 0.0793 00778 0.0764 0.0749 0.07350.072 0,0708 0.0694 0.0681 13 0.0968 0,0951 0.0934 0.0918 0.0901 0.0885 0.0869 00853 0.0838 0.0823 12 0.11503 0.1112 0.1093 0.1075 0.1056 01038 0,1020 01003 0.0985 -1 0.1357 0.1335 0.1314 0.1292 0.1271 0.125 0.1230 0.1210 0.1190 0.1170 10 0.15870.1562 0.1539 0.1515 0.1492 0.1469 0.1446 0.1423 0.1401 0.1379 -09 0.1841 0,1814 0.1788 0.1762 0.1736 0.1711 0.1685 0.1660 01635 0.161 0.8 0.21190.2090 0.206 0.2033 0.2005 0.19770.1949 0.1922 0.1894 0.1867 0.7 0,2420 0.2389 0.2358 0.2327 0.2296 0.2266 0.2236 0.2206 0.2177 0.2148 06 0.2743 0.27090.26760.2643 0.2611 0.2578 0.2546 0.2514 0.2483 0.2451 -0.5 0.3085 0.3050 0.3015 0.298 0.2946 0.2912 0.2877 0.2843 0.2810 0.2776 -04 0.34460.3409 03372 0.3336 0.3300 0.3264 0.3228 0.3192 0.3156 0.312 -0.3 0.3821 0.3783 03745 0.3707 0.3669 0.3632 03594 0.3557 0.3520 0.3483 -0.2 0.4207 0.4 168 0.4129 0.4090 0.4052 0.4013 0.3974 O.3936 0.3897 0.3859 -0.1 0.4602 0.4562 04522 0.4483 0.4443 0.4404 0.4364 0,4325 0.4286 0.4247 00 0.5000 0.4960 04920 04880 04840 0.48010.4761 0.4721 0.4681 0.4641 0.0548 0.0537 0052ABLE 1 CUMULATIVE PROBABILITIES FOR THE STANDARD NORMAL DISTRIBUTION (Continued) Cumulative probability Entries in the table give the area under the curve to the left of the z valuc. For example, for z-1.25, the cumulative probability is .8944 ,00 ,02 03 04 05 ,06 .07 .08 ,09 0 5000 5040 5080 5120 5160 5199 5239 5279 5319 5359 1 5398 5438 5478 5517 5557 5596 5636 5675 5714 5753 2 5793 5832 5871 5910 5948 5987 6026 6064 6103 6141 3 6179 6217 6255 6293 633 6368 6406 64436480 6517 4 6554 6591 6628 6664 6700 6736 6772 6808 6844 6879 5 6915 6950 6985 7019 7054 7088 7123 .7157 7190 7224 6 .7257 7291 7324 7357 7389 7422 7454 7486 .7517 .7549 7 .7580 7611 7642 7673 7704 7734 7764 7794 7823 7852 8 .7881 .7910 7939 .7967 7995 8023 8051 8078 8106 8133 .9 .8159 .8186 .8212 .8238 .8264 .8289 .8315 .8340 .8365 .8389 1.0 .8413 .8438 .8461 .8485 .8508 .8531 .8554 .8577 .8599 .8621 1.1 8643 8665 8686 8708 8729 8749 8770 8790 8810 8830 1.2 8849 8869 8888 8907 8925 8944 8962 8980 8997 9015 13 9032 9049 9066 9082 9099 9115 9131 9147 9162 9177 14 9192 9207 9222 9236 925 9265 9279 9292 9306 9319 15 9332 9345 9357 9370 9382 9394 9406 9418 9429 9441 1.6 9452 9463 9474 9484 9495 9505 9515 9525 9535 9545 1.7 9554 9564 9573 9582 9591 9599 9608 9616 9625 9633 18 9641 9649 9656 9664 9671 9678 9686 9693 9699 9706 1.9 9713 9719 9726 9732 9738 9744 9750 9756 9761 9767 2.0 9772 9778 9783 9788 9793 9798 9803 9808 .9812 9817 2.1 9821 9826 9830 9834 9838 9842 9846 9850 9854 .9857 2.2 9861 9864 9868 9871 9875 9878 9881 9884 .9887 .9890 2.3 9893 9896 9898 9901 9904 9906 9909 9911 9913 .9916 2.4 9918 9920 9922 9925 9927 9929 9931 9932 9934 9936 25 9938 9940 9941 9943 9945 9946 9948 9949 99519952 2.6 9953 9955 9956 9957 9959 9960 9961 9962 9963 9964 27 9965 9966 9967 9968 9969 9970 9971 9972 9973 9974 2.8 9974 9975 9976 9977 9977 9978 9979 9979 9980 9981 2.9 9981 9982 9982 9983 9984 9984 .9985 9985 9986 9986 3.0 9987 9987 9987 9988 9988 9989 9989 9989 .9990 9990eBook According to the Current Results website, the state of California has a mean annual rainfall of 21 inches, whereas the state of New York has a mean annual rainfall of 55 inches. Assume that the standard deviation for both states is 3 inches. A sample of 30 years of rainfall for California and a sample of 45 years of rainfall for New York has been taken. Use z-table. a. Show the probability distribution of the sample mean annual rainfall for California. This is a graph of a normal distribution with E( b. What is the probability that the sample mean is within 1 inch of the population mean for California? (to 4 decimals). 3 (to 4 decimals) c. What is the probability that the sample mean is within 1 inch of the population mean for New York? 3 (to 4 decimals) d. In which case, part (b) or part (c), is the probability of obtaining a sample mean within 1 inch of the population mean greater? Why? The probability of being within 1 inch is greater for New York in part (e) because the sample size is larger Hide Feedback Partially Correct
Answer #1

a) E(ar x) = 21

    sigma_{ar x} = sigma/sqrt n

          = 3/V30

          = 0.5477

b) P(20 < ar x < 22)

= P((20 - mu)/(sigma/sqrt n) < (ar x - mu)/(sigma/sqrt n) < (22 - mu)/(sigma/sqrt n))

= P((20 - 21)/0.5477 < Z < (22 - 21)/0.5477)

= P(-1.83 < Z < 1.83)

= P(Z < 1.83) - P(Z < -1.83)

= 0.9664 - 0.0336

= 0.9328

c) P(54 < ar x < 56)

= P((54 - mu)/(sigma/sqrt n) < (ar x - mu)/(sigma/sqrt n) < (56 - mu)/(sigma/sqrt n))

= P((54 - 55)/(3/V45) < Z < (56 - 55)/(3/V45))

= P(-2.24 < Z < 2.24)

= P(Z < 2.24) - P(Z < -2.24)

= 0.9875 - 0.0125

= 0.9750